Using = š›^0 H find the ratio E^0/H^0 for a plane electromagnetic wave propagating through vacuum. Show that it has the dimensions of electric resistance. This ratio is a universal constant called the impedance of free space.

Grade:10

1 Answers

Deepak Patra
9 years ago
Sol. E = Q/Ļµ base 0 A (Electric field) āˆ… = E.A. = Q/Ļµ base 0 A A/2 = Q/Ļµ base 0 2 i base 0 = Ļµ base 0 dāˆ… base E/dt = Ļµ base 0 d/dt (Q/Ļµ base 0 2) = 1/2 (dQ/dt)